Froyd Review & Training Center Inc. is expanding its school facilities starting 2001.  The program requires the following estimated expenditures: 

P1,000,000 at the end of 2001 

P1,200,000 at the end of 2002 

P1,500,000 at the end of 2003 

To accumulate the required funds, it establish a sinking fund constituting of 15 uniform  annual deposits, the first deposit has been made at the end of 1992. The interest rate of  the fund is 2% per annum. Calculate the annual deposit.



1
Expert's answer
2021-12-09T07:46:04-0500

Solution:

Estimated expenditures total = 1,000,000 + 1,200,000 + 1,500,000 = 3,700,000

"FV = P[\frac{(1+r)^{n}-1 }{r} ]"


"3,700,000 = P[\frac{(1+0.02)^{15}-1 }{0.02} ]"


P = 217,679.01

The annual deposit = 217,679.01